# Quantum Non-Local Nonstabilizerness

### v3

@@ Acknowledgements.
$P(y)=\sum_{x:\mathcal{M}(x)=y}\frac{P(x)}{\mathcal{M}^{\prime}(x)},$ (S10)
@@
$\mathcal{M}^{\prime}(x)=-\frac{4(2x-1)(1+8x(x-1))}{1+4(2x-1)^{2}(x-1)x}.$ (S11)
@@
$x=\frac{1}{2}\pm\frac{1}{4}\sqrt{2\pm 2e^{-y/2}z}.$ (S12)
For simplicity, we introduce the notation $z\equiv\sqrt{4-3e^{y}}$. Substituting Eq. (S11) and Eq. (S12) into Eq. (S10), we finally obtain the probability distribution of $y$ as:
$P(y)=\frac{3e^{-y/2}((e^{y/2}+z)\sqrt{1-e^{-y/2}z}+(e^{y/2}-z)\sqrt{1+e^{-y/2}z})}{4z\sqrt{2(e^{y}-1)}}.$ (S13)
@@
S1.4 Decay exponent for TFIM
@@
$\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})\sim r^{-\alpha_{\mu\nu}}+c_{\mu\nu},$ (S14)
@@
$M(\rho_{1,r})\sim\text{log}\left(\frac{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{2}}{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{4}}\right)\sim\text{log}\left(\frac{c_{1}+\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}}}{c_{2}+\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots}\right),$ (S15)
@@
$M(\rho_{1,r})\sim\text{log}\left(\frac{c_{1}}{c_{2}}\right)+\left[1+\frac{1}{c_{1}}(\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}})\right]\left[1-\frac{1}{c_{2}}(\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots)\right],$ (S16)
where $(1+x)^{-1}\sim(1-x)$ and $\text{log}(1+x)\sim x$ for small $x$ are used. Denoting $\alpha^{*}\equiv\min_{\mu\nu}\alpha_{\mu\nu}$ and $\overline{\alpha}\equiv\min_{\mu\nu}\{\alpha_{\mu\nu};c_{\mu\nu}\neq 0\}$, it follows that the in the large $r$ limit, the total nonstabilizerness $M$ is governed by:
$M(\rho_{1,r})\sim\text{const}+r^{-2\alpha^{*}}+r^{-\overline{\alpha}}.$ (S17)
@@
$\alpha_{M}=\text{min}(2\alpha^{*},\overline{\alpha}).$ (S18)
@@
At the critical point of the TFIM, we have $\alpha^{*}=\alpha_{xx}=1/4$ and $\overline{\alpha}=\alpha_{zz}=2$ [7], which results in $\alpha_{M}=1/2$, identical to the decay exponent of the mutual information. Furthermore, since the ground state is in canonical form, the NN is equivalent to the total nonstabilizerness of the two-qubit state. This completes the proof.
S2 Numerical results
S2.1 Numerical details
@@
In this section, we provide numerical details for the simulations conducted in the main text. The ground state of the TFIM is obtained using the density matrix renormalization group algorithm using the ITensor package [8, 9], with the maximum bond dimension $\chi=500$ and the truncation cutoff $\epsilon=10^{-12}$. For the monitored Haar-random unitary circuit at the critical point, the time evolution was performed by the time-evolving block decimation algorithm, with the maximum bond dimension $\chi=1000$ and the truncation cutoff $\epsilon=10^{-6}$.
@@
When calculating the NN via numerical optimization over all local unitary transformations, we parameterize the local unitary on a single qubit as $U=\begin{pmatrix}\text{cos}(\theta/2)&-e^{i\lambda}\text{sin}(\theta/2)\\ e^{i\phi}\text{sin}(\theta/2)&e^{i(\phi+\lambda)}\text{cos}(\theta/2)\end{pmatrix}$, where $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$ and $\lambda\in[0,2\pi]$. The optimization process employed the Nelder-Mead algorithm, as implemented in the Optim package [10]. To mitigate the risk of converging to a local minimum, we initiated the optimization from 100 distinct starting points and selected the minimal value achieved.
S2.2 Separable mixed states host nonzero NN
@@
In this section, we demonstrate that for any choice of nonstabilizerness measure for mixed states, certain separable states will always exhibit a nonzero NN. A separable state is defined as the convex hull of pure product states [11]:
$\rho=\sum_{i}p_{i}\rho_{iA}\otimes\rho_{iB}.$ (S19)
@@
Within the framework of resource theory, there are two inequivalent ways to define mixed stabilizer states, which we denote as $\text{STAB}_{0}$ and STAB, following the notation in Ref. [12]. $\text{STAB}_{0}$ is defined as the set of states $\rho=\frac{1}{2^{N}}\sum_{P\in G}P$, where $G$ is a group of commuting Pauli operators $P$ [13], while STAB is defined as the convex hull of pure stabilizer states:
$\text{STAB}:=\{\rho|\rho=\sum_{i}p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\},$ (S20)
@@
where $\psi_{i}$ are pure stabilizer states. When STAB is regarded as the set of free states, it specifically captures distillable non-Clifford resources. On the other hand, regarding $\text{STAB}_{0}$ as free states, it also incorporates the non-flat probability distributions of mixed states as resources. In the second SRE used in the main text, $\text{STAB}_{0}$ is treated as the set of free states, while other nonstabilizerness measures, such as the robustness of magic (RoM) [14, 15, 16] and mana [17], consider STAB as the set of free states. It is also worth mentioning that the SRE has been extended by a convex-roof construction to provide an alternative measure in the latter case [18].
@@
$\rho_{0}=\frac{1}{2}\left|\phi_{0}\phi_{0}\right\rangle\left\langle\phi_{0}\phi_{0}\right|+\frac{1}{2}\left|00\right\rangle\left\langle 00\right|.$ (S22)

### v4

@@ Appendix C Probability distribution
$P(y)=\sum_{x:\mathcal{M}(x)=y}\frac{P(x)}{\mathcal{M}^{\prime}(x)},$ (12)
@@
$\mathcal{M}^{\prime}(x)=-\frac{4(2x-1)(1+8x(x-1))}{1+4(2x-1)^{2}(x-1)x}.$ (13)
@@
$x=\frac{1}{2}\pm\frac{1}{4}\sqrt{2\pm 2e^{-y/2}z}.$ (14)
For simplicity, we introduce the notation $z\equiv\sqrt{4-3e^{y}}$. Substituting Eq. (13) and Eq. (14) into Eq. (12), we finally obtain the probability distribution of $y$ as:
$P(y)=$ $\frac{3e^{-y/2}}{4z\sqrt{2(e^{y}-1)}}\left[(e^{y/2}+z)\sqrt{1-e^{-y/2}z}\right.$ (15)
$+\left.(e^{y/2}-z)\sqrt{1+e^{-y/2}z}\right].$
@@
Appendix D Decay exponent for TFIM
@@
$\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})\sim r^{-\alpha_{\mu\nu}}+c_{\mu\nu},$ (16)
@@
$M(\rho_{1,r})\sim\text{ln}\left(\frac{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{2}}{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{4}}\right)$ (17)
$\sim\text{ln}\left(\frac{c_{1}+\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}}}{c_{2}+\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots}\right),$
@@
$M(\rho_{1,r})\sim\text{ln}\left(\frac{c_{1}}{c_{2}}\right)$ (18)
$+\left[1+\frac{1}{c_{1}}(\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}})\right]\times$
$\left[1-\frac{1}{c_{2}}(\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots)\right],$
where $(1+x)^{-1}\sim(1-x)$ and $\text{ln}(1+x)\sim x$ for small $x$ are used. Denoting $\alpha^{*}\equiv\min_{\mu\nu}\alpha_{\mu\nu}$ and $\overline{\alpha}\equiv\min_{\mu\nu}\{\alpha_{\mu\nu};c_{\mu\nu}\neq 0\}$, it follows that the in the large $r$ limit, the total nonstabilizerness $M$ is governed by:
$M(\rho_{1,r})\sim\text{const}+r^{-2\alpha^{*}}+r^{-\overline{\alpha}}.$ (19)
@@
$\alpha_{M}=\text{min}(2\alpha^{*},\overline{\alpha}).$ (20)
@@
At the critical point of the TFIM, we have $\alpha^{*}=\alpha_{xx}=1/4$ and $\overline{\alpha}=\alpha_{zz}=2$ Francesco et al. (2012), which results in $\alpha_{M}=1/2$, identical to the decay exponent of the mutual information. Furthermore, since the ground state is in canonical form, the NN is equivalent to the total nonstabilizerness of the two-qubit state. This completes the proof.
Appendix E Numerical details
@@
In this section, we provide numerical details for the simulations conducted in the main text. The ground state of the TFIM is obtained using the density matrix renormalization group algorithm using the ITensor package Fishman et al. (2022a, b), with the maximum bond dimension $\chi=500$ and the truncation cutoff $\epsilon=10^{-12}$. For the monitored Haar-random unitary circuit at the critical point, the time evolution was performed by the time-evolving block decimation algorithm, with the maximum bond dimension $\chi=1000$ and the truncation cutoff $\epsilon=10^{-6}$.
@@
When calculating the NN via numerical optimization over all local unitary transformations, we parameterize the local unitary on a single qubit as $U=\begin{pmatrix}\text{cos}(\theta/2)&-e^{i\lambda}\text{sin}(\theta/2)\\ e^{i\phi}\text{sin}(\theta/2)&e^{i(\phi+\lambda)}\text{cos}(\theta/2)\end{pmatrix}$, where $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$ and $\lambda\in[0,2\pi]$. The optimization process employed the Nelder-Mead algorithm, as implemented in the Optim package Mogensen and Riseth (2018). To mitigate the risk of converging to a local minimum, we initiated the optimization from 100 distinct starting points and selected the minimal value achieved.
Appendix F Nonzero NN in separable mixed states
@@
In this section, we demonstrate that for any choice of nonstabilizerness measure for mixed states, certain separable states will always exhibit a nonzero NN. A separable state is defined as the convex hull of pure product states Horodecki et al. (2009):
$\rho=\sum_{i}p_{i}\rho_{iA}\otimes\rho_{iB}.$ (21)
@@
Within the framework of resource theory, there are two inequivalent ways to define mixed stabilizer states, which we denote as $\text{STAB}_{0}$ and STAB, following the notation in Ref. Cao et al. (2024). $\text{STAB}_{0}$ is defined as the set of states $\rho=\frac{1}{2^{N}}\sum_{P\in G}P$, where $G$ is a group of commuting Pauli operators $P$ Nielsen and Chuang (2010), while STAB is defined as the convex hull of pure stabilizer states:
$\text{STAB}:=\{\rho|\rho=\sum_{i}p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\},$ (22)
@@
where $\psi_{i}$ are pure stabilizer states. When STAB is regarded as the set of free states, it specifically captures distillable non-Clifford resources. On the other hand, regarding $\text{STAB}_{0}$ as free states, it also incorporates the non-flat probability distributions of mixed states as resources. In the second SRE used in the main text, $\text{STAB}_{0}$ is treated as the set of free states, while other nonstabilizerness measures, such as the robustness of magic (RoM) Howard and Campbell (2017); Ahmadi et al. (2018); Heinrich and Gross (2019) and mana Veitch et al. (2014), consider STAB as the set of free states. It is also worth mentioning that the SRE has been extended by a convex-roof construction to provide an alternative measure in the latter case Leone and Bittel (2024).
@@
$\rho_{0}=\frac{1}{2}\left|\phi_{0}\phi_{0}\right\rangle\left\langle\phi_{0}\phi_{0}\right|+\frac{1}{2}\left|00\right\rangle\left\langle 00\right|.$ (24)